### v4

# The Value of Information in Human-AI Decision-making

Abstract
Multiple agents—including humans and AI models—are increasingly combined to make decisions with the expectation of achieving complementary performance, where the decisions they make together outperform those made individually. However, knowing how to improve the performance of collaborating agents is often difficult without knowing more about what particular information and strategies each agent employs. With a focus on human-AI pairings, we contribute a decision-theoretic framework for characterizing the value of information—and consequently, opportunities for agents to better exploit available information–in AI-assisted decision workflows. We present a novel explanation technique (ILIV-SHAP) that adapts SHAP explanations to highlight human-complementing information. We validate the effectiveness of the framework and ILIV-SHAP through a study of human-AI decision-making. We show that our measure of complementary information can be used to identify which AI model will best complement human decisions. We also find that presenting ILIV-SHAP with AI predictions leads to reliably greater reductions in error over non-AI assisted decisions more than vanilla SHAP.  [^1] The code to calculate the main quantities in our framework and reproduce the experimental results is available at https://osf.io/p2qzy/?view_only=bf39de5d96f047f69e45ffd42689ebf9.
1 Introduction
As the performance of artificial intelligence (AI) models continues to improve across domains, workflows in which human experts and AI models are paired for decision-making are sought in medicine, finance, and law, among others. Statistical models can often exceed the accuracy of human experts on average (Ægisdóttir et al., 2006; Grove et al., 2000; Meehl, 1954). However, whenever humans have access to additional information over an AI model, there is potential to achieve complementary performance by pairing the two, i.e., better performance than either the human or AI alone. For example, a physician may have access to information that is not captured in structured health records (Alur et al., 2024).
Evidence of complementary performance between humans and AI nonetheless remains limited, with many studies showing that human-AI teams underperform an AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019b; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001). A rigorous understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or agents individually using measures based on post-hoc decision accuracy (Passi and Vorvoreanu, 2022). However, such approaches do not account for the best achievable performance given the information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024). Moreover, existing approaches cannot provide insight on the potential for available information in the environment to improve decision performance. Consequently, it is difficult to design interventions that improve the team’s performance.
If we could identify information complementarities that can be exploited to maximize the decision performance of a human and AI model—such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated contextually-available information in their judgments—we can intervene to improve decision-making. For example, if we can identify how much complementary information over human judgments an AI model provides, we can use this knowledge to guide model selection or motivate further data collection to improve the model. Conversely, finding evidence that model predictions contain decision-relevant information that humans do not exploit can motivate the design of explanations highlighting complementary information.
We contribute a decision-theoretic framework for characterizing the value of information available within an AI-assisted decision workflow. In our framework, information is considered valuable to a decision-maker to the extent that it is possible, in theory, to incorporate it into their decisions to improve performance. Specifically, our approach analyzes the expected marginal payoff gain from best case (Bayes rational) use of additional information over best case use of the information already encoded in agent decisions. While we focus on combinations of humans and AI agents, the methods we propose can be applied to any combination of agents.
We introduce two metrics for evaluating information value in human-AI collaboration. The first—global human-complementary information value—calculates the value of a new piece of information to an agent over all of its possible realizations among all instances. The second—instance-level human-complementary information value—identifies opportunities for decision-makers to better use instance-level information such as specific AI predictions. We derive a new explanation technique (ILIV-SHAP) that reveals how data features influence the value of complementary information for an individual prediction.
To evaluate these tools, we contribute the results of a crowdsourced between-subjects experiment in which humans make decisions with and without AI models with varying human-complementary information and with access to different explanation approaches. We find that an AI model with more human-complementary information leads to greater improvements in human-AI team performance over the human-alone baseline. We also find that adding an ILIV-SHAP explanation to a traditional SHAP explanation leads to greater improvements in human-AI performance over the human-alone baseline than only the SHAP explanation or no explanation. We demonstrate further uses of the framework for model selection, empirical evaluation of human-AI performance, and explanation design on existing datasets spanning medical diagnosis, human deepfake detection, and recidivism prediction (see Appendices E, D and F). We also include an observational study on the dataset from Vodrahalli et al. (2022) which identified the AI model with more human-complementary information helps human-AI teams achieve greater improvements in performance over the human-alone baseline in Appendix C.
2 Related work
Human-AI complementarity.
Many empirical studies of human-AI collaboration focus on AI-assisted human decision-making for legal, ethical, or safety reasons (Bo et al., 2021; Boskemper et al., 2022; Bondi et al., 2022; Schemmer et al., 2022). However, a recent meta-analysis by Vaccaro et al. (2024) finds that, on average, human–AI teams perform worse than the better of the two agents alone. In response, a growing body of work seeks to evaluate and enhance complementarity in human–AI systems (Bansal et al., 2021b, 2019, a; Wilder et al., 2021; Rastogi et al., 2023; Mozannar et al., 2024b). The present work differs from much of this prior work by approaching human-AI complementarity from the perspective of information value and use, including asking whether the human and AI decisions provide additional information that is not used by the other.
Evaluation of human decision-making with machine learning.
Our work contributes methods for evaluating the decisions of human-AI teams (Kleinberg et al., 2015, 2018; Lakkaraju et al., 2017; Mullainathan and Obermeyer, 2022; Rambachan, 2024; Guo et al., 2024; Ben-Michael et al., 2024; Shreekumar, 2025). Kleinberg et al. (2015) proposed that evaluations of human-AI collaboration should be based on the information that is available at the time of decisions. According to this view, our work defines Bayesian best-attainable-performance benchmarks similar to several prior works (Guo et al., 2024; Wu et al., 2023; Agrawal et al., 2020; Fudenberg et al., 2022). Closest to our work, Guo et al. (2024) model the expected performance of a rational Bayesian agent faced with deciding between the human and AI recommendations as the theoretical upper bound on the expected performance of any human-AI team. This benchmark provides a basis for identifying exploitable information within a decision problem.
Human information in machine learning.
Some approaches focus on automating the decision pipeline by explicitly incorporating human expertise in developing machine learning models, such as by learning to defer (Mozannar et al., 2024a; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022, 2021; Okati et al., 2021). Corvelo Benz and Rodriguez (2023) propose multicalibration over human and AI model confidence information to guarantee the existence of an optimal monotonic decision rule. Alur et al. (2023) propose a hypothesis testing framework to evaluate the added value of human expertise over AI forecasts. Our work shares the motivation of incorporating human expertise, but targets a slightly broader scope by quantifying the information value for all available signals and agent decisions in a specific scenario of human–AI decision tasks.
3 Framework
Our framework takes as input a decision problem associated with an information model and outputs the value of information of any available signals to any agent, conditioning on the existing information in their decisions within a Bayesian decision theoretic framework. The intuition behind our use of statistical decision theory is that while human decision-makers may deviate from Bayesian rational decision-making, understanding the possible value of a signal requires defining the best attainable decision performance with that signal. Our framework provides two separate functions to quantify the value of information: one globally across the data-generating process, and one locally in a realization drawn from the data-generating process. We also introduce a robust analysis approach to information order, which enables us to compare the agent-complementary information in signals for all possible decision problems.
Decision Problem.
A decision problem consists of three key elements. We illustrate with an example of a weather decision.
  • A payoff-relevant state $\omega$ from a space $\mathbf{\Omega}$. For example, $\omega\in\mathbf{\Omega}=\{0,1\}=\{\text{no rain},\text{rain}\}$.
  • A decision $d$ from the decision space $\mathbf{D}$ characterizing the decision-maker (DM)’s choice. For example, $d\in\mathbf{D}=\{0,1\}=\{\text{not take umbrella},\text{take umbrella}\}$.
  • A payoff function $S:\mathbf{D}\times\mathbf{\Omega}\to\mathbb{R}$, used to assess the quality of a decision given a realization of the state. For example, $S(d=0,\omega=0)=0,S(d=0,\omega=1)=-100,S(d=1,\omega=0)=-50,S(d=1,\omega=1)=0$, which punishes the DM for selecting an action that does not match the weather.
In decision problems corresponding to prediction tasks, the decision space is a probabilistic belief over the state space, i.e., $\mathbf{D}=\Delta(\mathbf{\Omega})$. For such problems, a payoff function is said to be a proper scoring rule if the optimal action is to predict the true distribution, i.e., $p=\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim p}\![{S(d,\omega)}]$. For any decision problem with payoff function $S:\mathbf{D}\times\mathbf{\Omega}\to\mathbb{R}$, there is an equivalent proper scoring rule $\hat{S}:\Delta(\mathbf{\Omega})\times\mathbf{\Omega}\to\mathbb{R}$ defined by choosing the optimal decision under the reported belief. Formally,
$\hat{S}(p,\omega)=S(\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim p}\![{S(d,\omega)}],\omega).$ (1)
Equation 1 shows a reduction from the payoff function $S$ to the proper scoring rule $\hat{S}$, i.e., any decision $p$ under $\hat{S}$ represents a decision $d$ in $S$ that best-responds to the distribution $p$. Therefore, the best-attainable performance defined in the proper scoring rule $\hat{S}$ is equivalent to the best-attainable performance defined in any payoff function $S$ that can be reduced to $\hat{S}$. Throughout our framework, we prefer proper scoring rules over non-proper scoring rules, since the best-attainable performance defined in the former implies the best-attainable performance in the latter.
Information Model.
We cast the information available to a DM as a signal defined within an information model. We use the definition of an information model in Blackwell et al. (1951). The information model can be represented by a data-generating model with a set of signals.
  • Signals. There are $n$ “basic signals” represented as random variables $\Sigma_{1},\ldots,\Sigma_{n}$, from the signal spaces $\mathbf{\Sigma}_{1},\ldots,\mathbf{\Sigma}_{n}$. Basic signals represent information available to a decision-maker, e.g., $\mathbf{\Sigma}_{1}=\{\text{cloudy},\text{not cloudy}\}$, $\mathbf{\Sigma}_{2}\in\{0,\ldots,100\}$ for temperature in Celsius, etc. The decision-maker observes a signal, which is a subset of the basic signals, $V\subseteq 2^{\{\Sigma_{1},\dots,\Sigma_{n}\}}$. Specifically, we use $V=\{\Sigma_{j_{1}},\ldots,\Sigma_{j_{k}}\}$ for a signal having $k$ basic signals and denote the signal space as $\mathbf{V}=\mathbf{\Sigma}_{j_{1}}\times\ldots\times\mathbf{\Sigma}_{j_{k}}$. For example, a signal representing a combination of two basic signals $V=\{\Sigma_{1},\Sigma_{2}\}$ observed by the decision-maker might consist of cloudiness $\Sigma_{1}$ and the temperature $\Sigma_{2}$ of the day. Given a signal composed of $m$ basic signals, we write the realization of $V$ as $v=(\sigma_{j_{1}},\dots,\sigma_{j_{k}})$, where the realizations $\sigma_{j_{i}}\in\mathbf{\Sigma}_{j_{i}}$ are sorted by the index of the basic signals $j_{i}\in[n]$. The union $V$ of two signals $V_{1},V_{2}$ takes the set union, i.e., $V=V_{1}\cup V_{2}$.
  • Data-generating process. A data-generating process is a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega})$ over the basic signals and the payoff-relevant state. $\pi$ can be viewed as the combination of two distributions: the prior distribution of the state $\Pr[\omega]$ and the signal-generating distribution $\Pr[v|\omega]$ defining the conditional distribution of signals. The DM may only observe a subset $V$ of the $n$ basic signals. We can define the Bayesian posterior belief upon receiving a signal $V=v$ from the data-generating model as
    $\pi(\omega|v):=\Pr[\omega|v]=\frac{\pi(v,\omega)}{\pi(\omega)}$
    where we slightly abuse notation to write $\pi(v,\omega)$ as the marginal probability of the signal realized to be $v$ and the state being $\omega$ with expectation over other signals.
The choice of basic signals directly impacts how many observed samples are required to get a good estimate of the data-generating process. When high-dimensional signals such as images or text are used, it may not be computationally feasible to estimate the data-generating process. In such cases, one can preprocess the high-dimensional signals to get low-dimensional representations which (ideally) capture any important structure. These lower-dimensional signals can be defined by humans, such as when concepts are identified and then used to label the high-dimensional signals (e.g., images or parts of images). Alternatively, they can be defined algorithmically, by strategically applying dimensionality reduction to generate low-dimensional embeddings. We introduce an algorithm for identifying a set of optimal “coarsened” signals given a set of observed samples in Appendix B.
Information value.
Our framework quantifies the value of information in a signal $V$ as the expected payoff improvement of an idealized agent who has access to $V$ in addition to some baseline information set. This corresponds to a rational Bayesian DM who knows the prior probability of the state and conditional distribution of signals (i.e., the data-generating process), observes a signal realization, updates their prior to arrive at posterior beliefs, and then chooses a decision to maximize their expected payoff given their posterior belief. Formally, given a decision task with payoff function $S$ and an information model $\pi$, the rational DM’s expected payoff given a (set of) signal(s) $V$ is
$\mathrm{R}^{\pi,S}(V)=\text{\bf E}_{(v,\omega)\sim\pi}\![{S(d^{r}(v),\omega)}]$ (2)
where $d^{r}(\cdot):\mathbf{V}\rightarrow\mathbf{D}$ denotes the decision rule adopted by the rational DM.
$d^{r}(v)=\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d,\omega)}]$ (3)
We use $\emptyset$ to represent a null signal, such that $\mathrm{R}^{\pi,S}(\emptyset)$ is the expected payoff of a Bayesian rational DM who only uses the prior distribution to make decisions. In this case, the Bayesian rational DM will take the best fixed action under the prior, and their expected payoff is:
$\mathrm{R}^{\pi,S}(\emptyset)=\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi}\![{S(d,\omega)}]$ (4)
This baseline defines the maximum expected payoff that can be achieved with no information. Bayesian decision theory quantifies the information value of $V$ as the payoff improvement of $V$ over the payoff obtained without information.
Definition 3.1.
Given a decision task with payoff function $S$ and an information model $\pi$, the information value of $V$ is defined as
$\operatorname{IV}^{\pi,S}(V)=\mathrm{R}^{\pi,S}(V)-\mathrm{R}^{\pi,S}(\emptyset)$ (5)
We adopt the same idea to define the agent-complementary information value in our framework.
3.1 Agent-Complementary Information Value
With the above definitions, it is possible to measure the best case additional value that new signals can provide over the information already captured by an agent’s decisions. Here, agent may refer to a human, an AI system, or a human–AI team. The intuition behind our approach is that any information that is used by decision-makers should eventually reveal itself through variation in their decisions. We recover the information value in agent decisions by offering the decisions as a signal to the Bayesian rational DM. We model the agent decisions as a random variable $D^{b}$ from the action space $\mathbf{D}$, which follows a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega}\times\mathbf{D})$ with the state and signals. The expected payoff of a Bayesian rational DM who knows $\pi$ is given by the function:
$\mathrm{R}^{\pi,S}(D^{b})=\text{\bf E}_{(d^{b},\omega)\sim\pi}\![{S(d^{r}(d^{b}),\omega)}]$
We seek to identify signals $V$ that can potentially improve agent decisions by analyzing the information value in the combined signal $D^{b}\cup V$ over the information value in $D^{b}$, which we define as the agent-complementary information value.
Definition 3.2.
Given a decision task with payoff function $S$ and an information model $\pi$, we define the agent-complementary information value ($\operatorname{ACIV}$) of $V$ on agent decisions $D^{b}$ as
$\operatorname{ACIV}^{\pi,S}(V;D^{b})=\mathrm{R}^{\pi,S}(D^{b}\cup V)-\mathrm{R}^{\pi,S}(D^{b})$ (6)
If the $\operatorname{ACIV}$ of a signal $V$ is small relative to the baseline (4), this means either that the information value of $V$ to the decision problem is low (e.g., it is not correlated with $\omega$), or that the agent has already exploited the information in $V$ (e.g., the agent relies on $V$ or equivalent information to make their decisions such that their decisions correlate with $\omega$ in the same way as $V$ correlates with $\omega$). If, however, the $\operatorname{ACIV}$ of $V$ is large relative to the baseline, then at least in theory, the agent can improve their payoff by incorporating $V$ in their decision making.
Furthermore, $\operatorname{ACIV}$ can reveal complementary information between different types of agents. For instance, if we view AI predictions as $V$ and treat human decisions as the agent signal $D^{b}$, a large $\operatorname{ACIV}$ indicates that AI predictions add considerable value beyond what humans alone achieve. In the reverse scenario, if human decisions serve as $V$ and AI predictions are $D^{b}$, we can measure how much humans can contribute over the information captured in the AI predictions. We demonstrate uses of $\operatorname{ACIV}$ in Appendix D and Appendix E.
3.2 Instance-level Agent-complementary Information Value
$\operatorname{ACIV}$ quantifies the value of the decision-relevant information in a signal $V$ across all possible realizations defined by the data-generating model. To provide analogous instance-level quantification of information value, we define Instance-Level agent-complementary Information Value ($\operatorname{ILIV}$) to quantify the value of the decision-relevant information encoded by a single realization of the signal. This finer-grained view makes it possible to analyze how much an agent can benefit in theory from better incorporating instance-level information in their decision.
Given a realization of signal $v=\{\sigma_{j_{1}},\ldots,\sigma_{j_{k}}\}$, we want to know the maximum expected payoff gain from gaining access to $v$ on the instances where $v$ is realized over the existing information encoded in agent decisions. Intuitively, this captures how much “room” there is for specific decisions to be improved through better use of the signal. To calculate instance-level information value, we rely on the performance of the Bayesian rational agent on the conditional distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b}|v)$ instead of the whole distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b})$. Formally, given a decision task with payoff function $S$ and information model $\pi$, the expected payoff of the rational DM given signal $V=v^{\prime}$ on instances with the signal realization $V=v$ is
$\mathrm{r}^{v,\pi,S}(v^{\prime})=\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d^{r}(v^{\prime}),\omega)}]$ (7)
where $d^{r}(v)$ is the Bayesian optimal decision on receiving $v^{\prime}$ as defined in Equation 3. If we consider the agent decisions in addition to the realization $v$, the rational DM’s expected payoff on instances where $V=v$ can be written as
$\mathrm{r}^{v,\pi,S}(v^{\prime};D^{b})=\text{\bf E}_{(d^{b},\omega)\sim\pi(d^{b},\omega|v)}\![{S(d^{r}(v^{\prime}\cup d^{b}),\omega)}]$ (8)
Note that we allow $v^{\prime}\neq v$, i.e. the signal $v^{\prime}$ observed by the rational DM being different from the actual realization $v$. When $v^{\prime}\neq v$, the rational DM is misinformed, thus $\mathrm{r}^{v,\pi,S}(v^{\prime})\leq\mathrm{r}^{v,\pi,S}(v)$. $\operatorname{ILIV}$ in Definition 3.3 quantifies this improvement in payoff when the rational DM observes the actual $v$.
Definition 3.3.
Given a decision task with payoff function $S$ and an information model $\pi$, we define the instance-level agent-complementary information value ($\operatorname{ILIV}$) of signal $V=v^{\prime}$ on instances where $V=v$ as:
$\operatorname{ILIV}^{v,\pi,S}(v^{\prime};D^{b})=\mathrm{r}^{v,\pi,S}(v^{\prime};D^{b})-\mathrm{r}^{v,\pi,S}(\emptyset;D^{b}).$ (9)
$\mathrm{r}^{v,\pi,S}(\emptyset;D^{b})$ represents the expected rational payoff on instances with $V=v$, where the rational DM only knows the information encoded in agent decisions $D^{b}$ without knowing any realizations of $V$. $\operatorname{ILIV}$ maximizes when the signal does not misinform, i.e., $\operatorname{ILIV}^{v,\pi,S}(v;D^{b})\geq\operatorname{ILIV}^{v,\pi,S}(v^{\prime};D^{b}),\text{for }v^{\prime}\in\mathbf{V}$.
3.2.1 ILIV-SHAP Information-based Explanation
We apply $\operatorname{ILIV}$ to define an information-based explanation technique (ILIV-SHAP) that extends the canonical SHAP feature saliency explanation of a model’s prediction. ILIV-SHAP communicates how the data features lead to changes in the information value of AI predictions. While traditional SHAP summarizes the average contribution of each feature to a specific prediction over the baseline (prior) prediction, ILIV-SHAP summarizes the average contribution of each feature to the decision-relevant information value contained in the prediction.
Suppose a model $f$ that takes as input $m$ features and outputs a real number. Given an instance $\mathbf{x}=(x_{1},\ldots,x_{m})$, the importance of one feature $x_{i}$ to the model output $f(\mathbf{x})$ is encoded by the expected difference of model outputs when $x_{i}$ is marginalized out. Specifically, this is quantified by $f(\mathbf{x})-\text{\bf E}\![{f(X)|X_{-i}=\mathbf{x}_{-i}}]$, where $X_{-i}$ denotes all features except $X_{i}$. Considering the interaction between features, SHAP (Lundberg and Lee, 2017) uses the Shapley value to quantify the importance scores averaged on different combinations of features:
$\phi_{i}(f,\mathbf{x})=\sum_{\mathbf{x}^{\prime}\subseteq\mathbf{x}}\frac{|\mathbf{x}^{\prime}|!(m-|\mathbf{x}^{\prime}|-1)!}{m!}[g_{f}(\mathbf{x}^{\prime})-g_{f}(\mathbf{x}^{\prime}\backslash x_{i})]$
where $g_{f}(\mathbf{x}^{\prime})$ denotes the expected model output conditioned on $\mathbf{x}^{\prime}$, i.e., $\text{\bf E}\![{f(X)|X^{\prime}=\mathbf{x}^{\prime}}]$ for any $\mathbf{x}^{\prime}\subseteq\mathbf{x}$. The scores $\phi_{i}(f,\mathbf{x})$ output by SHAP construct an explanation model for a model output, which quantifies the expected counterfactual change in the model output caused by the feature $x_{i}$.
ILIV-SHAP extends SHAP to give an explanation model of how features impact the decision-relevant information value of an individual model output.
Definition 3.4 (ILIV-SHAP).
Given a model $f$ and data features $\mathbf{x}=(x_{1},\ldots,x_{m})$, the importance score of the i-th feature by ILIV-SHAP is
$\phi_{i}^{\operatorname{ILIV}}(f,\mathbf{x})=\sum_{\mathbf{x}^{\prime}\subseteq\mathbf{x}}\frac{|\mathbf{x}^{\prime}|!(m-|\mathbf{x}^{\prime}|-1)!}{m!}[\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})]$
where $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})$ denotes a counterfactual evaluation of $\operatorname{ILIV}$, which quantifies the expected payoff gain from additionally knowing $g_{f}(\mathbf{x}^{\prime})$ on the instances where $f(\mathbf{x})$ is realized. This counterfactual version of $\operatorname{ILIV}$ is guaranteed to achieve the maximum when $g_{f}(\mathbf{x}^{\prime})$ informs $f(\mathbf{x})$, i.e., $f(\mathbf{x})=g_{f}(\mathbf{x}^{\prime})$ – the features missing from $\mathbf{x}^{\prime}$ have no impact on $f(\mathbf{x})$.
The explanation model offered by ILIV-SHAP is grounded in the following two properties. First, $\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$ are consistent with the extent to which feature $x_{i}$ contributes to the information value in $f(\mathbf{x})$. Specifically, for any two models $f$ and $f^{\prime}$ and any $\mathbf{x}^{\prime}\subseteq\mathbf{x}$, if $x_{i}$ contributes more to the information value in $f^{\prime}$ than the information value in $f$, i.e., $\operatorname{ILIV}^{f^{\prime}(\mathbf{x}),\pi,S}(g_{f^{\prime}}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f^{\prime}(\mathbf{x}),\pi,S}(g_{f^{\prime}}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})\geq\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})$, then $\phi^{\operatorname{ILIV}}_{i}(f^{\prime},\mathbf{x})\geq\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$. This property of ILIV-SHAP follows the consistency property of SHAP (Lundberg and Lee, 2017). Second, summing up the scores $\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$ recovers the information value of model output, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})$. Formally, for any $f$ and any $\mathbf{x}$, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})=\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(\text{\bf E}\![{f(X)}];D^{b})+\sum_{i=1}^{m}\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$. We demonstrate use of ILIV-SHAP in Appendix F.
3.3 Robust Analysis of Information Order
Our approach assumes a decision problem as input and evaluates agents’ decisions and use of information on this problem. However, evaluators may face ambiguity around the appropriate decision problem specification, and in particular, the appropriate scoring rule. In particular, ambiguity can arise in payoff functions; doctors, for example, penalize false negative results differently when diagnosing younger versus older patients (Mclaughlin and Spiess, 2023). Blackwell’s comparison of signals (Blackwell et al., 1951) is an ideal tool for addressing ambiguity about the payoff function, as it defines a signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems. We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules [^2] For rational DMs, any decision problem can be represented by an equivalent proper scoring rule in Equation 1, such that the partial order defined via proper scoring rules also applies to the corresponding decisin tasks. (Li et al., 2022; Kleinberg et al., 2023).
Definition 3.5 (Blackwell Order of Information).
A signal $V_{1}$ is Blackwell more informative than $V_{2}$ if $V_{1}$ achieves a higher best-attainable payoff on any decision problems:
$\mathrm{R}^{\pi,S}(V_{1})\geq\mathrm{R}^{\pi,S}(V_{2}),\forall S$
The Blackwell order is evaluated over all possible decision problems, which cannot be tested directly. Fortunately, we only need to test over all proper scoring rules since any decision problem can be represented by an equivalent proper scoring rule, and the space of proper scoring rules can be characterized by a set of V-shaped scoring rules. A V-shaped scoring rule is parameterized by the kink of the piecewise-linear utility function.
Definition 3.6.
(V-shaped scoring rule) A V-shaped scoring rule with kink $\mu\in(0,\frac{1}{2}]$ is defined as
$S_{\mu}(d,\omega)=\left\{\begin{array}[]{cc}\frac{1}{2}-\frac{1}{2}\cdot\frac{\omega-\mu}{1-\mu}&\text{if }d\leq\mu\\ \frac{1}{2}+\frac{1}{2}\cdot\frac{\omega-\mu}{1-\mu}&\text{else},\end{array}\right.$
When $\mu^{\prime}\in(\frac{1}{2},1)$, the V-shaped scoring rule can be symmetrically defined by $S_{\mu^{\prime}}=S_{1-\mu^{\prime}}(1-y,\omega)$.
Intuitively, the kink $\mu$ represents the threshold belief where the decision-maker switches between two actions. Larger $\mu$ means that the decision-makers will prefer $d=1$ more. The closer $\mu$ is to $0.5$, the more indifferent the decision-maker is to $d=0$ or $d=1$.
Proposition 3.7 shows that if $V_{1}$ achieves a higher information value on the basis of V-shaped proper scoring rules than $V_{2}$, then $V_{1}$ is Blackwell more informative than $V_{2}$. Proposition 3.7 follows from the fact that any best-responding payoff can be linearly decomposed into the payoff on V-shaped scoring rules.
Proposition 3.7 (Hu and Wu 2024).
If $\forall\mu\in(0,1)$
$\mathrm{R}^{\pi,S_{\mu}}(V_{1})\geq\mathrm{R}^{\pi,S_{\mu}}(V_{2}),$
then $V_{1}$ is Blackwell more informative than $V_{2}$.
Extending this to ACIV, $V_{1}$ offers a higher complementary value than $V_{2}$ under the Blackwell order if
$\operatorname{ACIV}^{\pi,S_{\mu}}(V_{1};D^{b})\geq\operatorname{ACIV}^{\pi,S_{\mu}}(V_{2};D^{b}),\forall\mu\in(0,1)$
This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Appendix D.
4 Demonstrations
We demonstrate uses of our framework on three real-world decision-making tasks: medical diagnosis, deepfake video detection, and recidivism prediction.
First, we demonstrate the framework’s utility in model selection by evaluating how well different cardiac dysfunction prediction models trained on chest X-rays complement radiologists’ predictions (Rajpurkar et al., 2018; Johnson et al., 2019), showing how even among models with similar accuracy,some models strictly offer more complementary information than others over the radiologists. Next, we show how the framework can be used to evaluate how well humans exploit available features in a deepfake detection task (Dolhansky et al., 2020; Groh et al., 2022), how this compares to how an AI models exploits the information, and how providing AI assistance affects the human use. Suprisingly, even though AI assistance helps human-AI teams achieve higher decision accuracy, the $\operatorname{ACIV}$s of the video-level features did not change much compared to the $\operatorname{ACIV}$s over human-only predictions, indicating that simply displaying AI’s prediction did not help humans exploit the available information. Lastly, we demonstrate how ILIV-SHAP explanations surface unique insight into information value that humans lack relative to conventional SHAP for a recidivism prediction task (Angwin et al., 2022; Dressel and Farid, 2018). Detailed descriptions and results of all demonstrations are in Appendices E, D and F.
5 Experiment
We use a preregistered between-subjects online experiment to answer two questions:
  1. 1.
    Can ACIV identify which AI model will result in the best human-AI decision-making?
  2. 2.
    Does ILIV-SHAP improved human-AI decision-making over SHAP alone?
Our experiment assigns participants to one of two AI models with varying ex-ante ACIV and one of three explanation conditions: ILIV-SHAP and SHAP, SHAP, and no explanation. We hypothesize that the higher ACIV model will lead to better human-AI decisions and presenting ILIV-SHAP explanations for this model will lead to better human-AI decisions than the other explanation types.
Metric AI1 AI2
MAPE 14.30% 14.51%
R2 0.81 0.81
$\operatorname{ACIV}$
(in MSE)
$6.5\times 10^{8}$ $3.7\times 10^{8}$
$\operatorname{ACIV}$
(in MAPE)
4.61% 2.00%
Figure 1: Left: AI performance and ACIV. Right: Posterior predicted reduction in the human-AI team’s absolute percentage error (APE) with 95% confidence intervals.
Task and data.
We study an AI-assisted house price prediction, following other recent crowdsourced studies (e.g., Chiang and Yin (2021); Hemmer et al. (2022); Poursabzi-Sangdeh et al. (2021); Holstein et al. (2023)) . This task does not require any domain knowledge, making it suitable for a broad participant population. Additionally, the task is complex enough to benefit from AI assistance. We use the Ames, Iowa Housing Dataset (De Cock, 2011).
Participants and procedure.
We recruited 421 US-based participants via the crowdsourcing platform Prolific. Each participant was randomly assigned to 12 houses in the Ames Iowa Housing Dataset and performed a sequence of 24 trials (where two decisions were elicited for each house, one without AI assistance and one with), split evenly across a human-alone prediction phase and a human-AI prediction phase. In the first 12 trials, participants were asked to predict the house price of the assigned house using six features of the house, including year built, the living area above ground in square feet, the size of garage in car capacity, the number of fireplaces, year remodeled, and ratings of the house’s material and finish. We intentionally reduced the interpretability of the year remodeled and material and finish by witholding what variables they represented and rescaling their values. In the second 12 trials, participants were asked to revise their guesses in the first round after seeing the AI’s prediction and explanation. We incentivized participants with a base reward of $3.00$ USD and a bonus based on the mean squared error (MSE) between their guesses and the true house price in the 24 trials: $\textit{bonus}=3.00-\frac{\textit{MSE}}{3\times 10^{9}}$.
AI models and explanations.
Participants were assigned to one of six experimental conditions resulting from crossing the AI models (2 levels, where AI1 has high $\operatorname{ACIV}$ and AI2 has low $\operatorname{ACIV}$) with the explanation type (3 levels: ILIV-SHAP + SHAP, SHAP, and no explanation). We designed AI1 and AI2 so as to vary their potential to complement humans: AI1 was trained with the six features and the true house price, with noise added to the six features. On the other hand, AI2 was trained with the four human-interpretable features and the true house price, such that there would be less potential for the predictions to add unique information over human judgments. We used the same training data for both AI models and ensured that the two AI models achieved similar performance (Figure 1), so that model accuracy differences would not confound our results. In the SHAP explanation, we used the SHAP values of the six features to explain the AI’s prediction. To generate the ILIV-SHAP explanations, we used independent human decisions we collected on Prolific for the dataset using the same set-up prior to running the official study. We present ILIV-SHAP explanations of the six features that ranked and highlighted the features where the human-complementary information the AI’s prediction gets from the feature is higher than a threshold (see Appendix I for explanation examples).
Evaluation metrics.
We report the $\operatorname{ACIV}$ of the AI models in the mean squared error (MSE) [^3] We included MSE since it is a proper scoring rule while MAPE is not. and the mean absolute percentage error (MAPE). To measure the human unassisted and human-AI team performance, we fit a preregistered Bayesian hierarchical regression model [^4] https://aspredicted.org/jfm6-rrs3.pdf to the percentage error (PE) between the human’s prediction and the true house price, $\textit{PE}=(d-\omega)/\omega$ using R’s brms package (Bürkner, 2017).
[…]
$\nu$ $\sim\text{Gamma}(2,0.1)$
[…]
Results.
Figure 1 confirms the expected ranking of AI models by $\operatorname{ACIV}$. AI1, which had access to the two human-uninterpretable features, achieves higher $\operatorname{ACIV}$ than AI2, which only had access to the four human-interpretable features. Note the similar predictive performance of the two AI models.
Posterior predictions of human-AI team performance according to our regression model also align with our expectations based on the complementary information each model provides (Figure 1 far right). Collapsing across explanation conditions, AI1 reduces APE by more than AI2 ($6.27\%[5.91\%,6.63\%]$ vs $5.65\%[5.26\%,6.05\%]$ respectively). Access to the ILIV-SHAP + SHAP explanation results in a greater reduction in APE for AI1 than AI2 ($7.11\%[6.57\%,7.65\%]$ vs $5.83\%[5.17\%,6.49\%]$ respectively).
Figure 1 also illustrates the effect of the different explanation conditions on the human-AI team performance. When the AI has sufficient complementing information (i.e., AI1), the ILIV-SHAP + SHAP explanation more effectively reduces APE for the human-AI decisions than the SHAP explanation and baseline (no explanation) ($7.11\%[6.57\%,7.65\%]$ for ILIV-SHAP, $5.20\%[4.53\%,5.87\%]$ for SHAP, and $6.50\%[5.86\%,7.14\%]$ for no explanation, all for model AI1).
6 Discussion and Limitations
Our decision-theoretic framework quantifies the information value of signals available in a human-AI decision setting over the information value of agent decisions. The basis of our framework in Bayesian decision theory means that it provides a theoretical basis that can be adapted to support comparisons to human behavior to drive learning (Hullman and Gelman, 2021), even if human agents deviate considerably from Bayesian rational decision-making. We experimentally demonstrate the power of the framework for analyzing and improving human-AI decision-making. Our proposed ILIV-SHAP explanation improves performance over an existing state-of-the-art explanation strategy. Our work provides a promising methodological framework for an emerging research agenda around optimally combining agents’ information for decisions (e.g., Alur et al. (2023, 2024)). While we focused on human-AI decisions, the framework an be applied to any combination of AI or human agent judgments.
One direction for follow-up work is to develop tests that complement the framework by helping evaluators confirm the information that agents are using. Note that low ACIV of a signal over human decisions is not necessarily sufficient for concluding that particular signals were used by the human. It is often unknown whether a human has other private signals offering equivalent information to a feature being analyzed.
Appendix A The Combinatorial Nature of the Value of Signals
When decision-makers are provided with multiple signals, the signals have the combinatorial property by nature. Acknowledgd by recent works in decision theory and game theory [Chen and Waggoner, 2016], one signal may have no information value by itself, but it can be complementary to other signals to provide information value. For example, two signals $\Sigma_{1}$ and $\Sigma_{2}$ are uniformly random bits and the state $\omega=\Sigma_{1}\oplus\Sigma_{2}$, the XOR of $\Sigma_{1}$ and $\Sigma_{2}$. In this case, neither of the signals offers information value on its own, but knowing both leads to the maximum payoff. Though we did not explicitly observe the complementation between signals in our survey of human-AI decision-making tasks (see the results of deepfake detection in supplementary materials), we want to note that our framework can be extended to consider this complementation between signals. We use the Shapley value [Shapley, 1953] to interpret the contribution to the $\operatorname{ACIV}$ of each basic signal. $\phi$ is the average of the marginal contribution of a signal in every combination with other signals.
$\phi(V)=\frac{1}{n}\sum_{V^{\prime}\subseteq\{\Sigma_{1},\ldots,\Sigma_{n}\}/V}{(n-\lvert V\rvert)\choose\lvert V^{\prime}\rvert}^{-1}(\operatorname{ACIV}^{\pi,S}(V^{\prime}\cup V;D^{b})-\operatorname{ACIV}^{\pi,S}(V^{\prime};D^{b}))$ (10)
The following algorithm provides a polynomial-time approximation of the Shapley value of $\operatorname{ACIV}$. Under the assumption of submodularity, it orders the signals the same as the Shapley value.
Algorithm 1 Greedy algorithm for marginal gain of $\operatorname{ACIV}$
1:  $V^{*}=\{D^{b}\}$
2:  $\Phi^{*}=\{\}$
3:  for $i=1$ to $n$ do
4:     $\phi^{\prime}_{j}=\operatorname{ACIV}(\Sigma_{j};V^{*})\ for\ each\ j$
5:     $j^{*}=\arg\max_{j\ s.t.\ \Sigma_{j}\notin V^{*}}\phi^{\prime}_{j}$
6:     $\phi_{j^{*}}=\max_{j\ s.t.\ \Sigma_{j}\notin V^{*}}\phi^{\prime}_{j}$
7:     $add\ \Sigma_{j^{*}}\ to\ V^{*}$
8:     $add\ \phi_{j^{*}}\ to\ \Phi^{*}$
9:  end for
10:  $output\ \phi_{j^{*}}$
Appendix B Estimating the Data-Generating Process from Observations
Suppose the data-generating process is estimated from the empirical distribution of the observed realizations of the payoff state value and the signal, $D=\{(\omega_{i},v_{i})\}_{i=1}^{T}$, i.e.,
$\hat{\pi}_{D}(\omega,v)=\frac{1}{T}\sum_{i=1}^{T}\mathbb{1}_{\omega=\omega_{i}}\cdot\mathbb{1}_{v=v_{i}}$ (11)
In applying our framework, the rational DM’s decision rule $d^{r}$ should not overfit to the observed dataset. This threatens the usefulness of analysis using the DGP. To avoid this situation, we can either collect more observations or “coarsen” the signal space, where “coarsen” means reducing the signal space by aggregating similar signals into a single signal. We propose a systematic procedure for the latter since it may be too costly or otherwise not possible to collect more observations. Our objective is to find the coarsened signal structure that has the largest signal space and avoids the rational decision rule from overlapping. Denote the maximum tolerance factor for overfitting as $\epsilon$. The rational DM’s decision rule $d^{r}_{\hat{\pi}_{D^{\prime}}}$ that is estimated from a set of observations $D^{\prime}$ should perform similarly on $D^{\prime}$ and on the evaluation of a holdout set $D\setminus D^{\prime}$, i.e.,
$\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]-\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D\setminus D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]\leq\epsilon*\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]$ (12)
Algorithm 2 Estimating the data-generating process from a set of observations
1:  Input: Observed dataset $C=\{(\omega_{i},v_{i})\}_{i=1}^{T}$, Overfitting tolerance $\epsilon$, Ratio of the holdout set $\delta$.
2:  randomly split the dataset into a training set $C^{\prime}$ and a holdout set $C\setminus C^{\prime}$ with
$\lvert C^{\prime}\rvert=(1-\delta)*T$
3:  $\hat{\pi}_{C^{\prime}}(\omega,v)=\frac{1}{T}\sum_{\hat{\omega},\hat{v}\in C^{\prime}}\mathbb{1}_{\omega=\hat{\omega}}\cdot\mathbb{1}_{v=\hat{v}}$
4:  $R_{training}=\text{\bf E}_{(v,\omega)\sim C^{\prime}}\![{S(d^{r}_{\hat{\pi}_{C^{\prime}}}(v),\omega)}]$.
5:  $R_{testing}=\text{\bf E}_{(v,\omega)\sim C\setminus C^{\prime}}\![{S(d^{r}_{\hat{\pi}_{C^{\prime}}}(v),\omega)}]$.
6:  $bin\_size=\text{the number of unique }v$
7:  if $R_{training}-R_{testing}>\epsilon*R_{training}$ then
8:     $bin\_size=bin\_size/2$
9:     $\{v^{\prime}_{i}\}_{i=1}^{T}=\text{coarsen }\{v_{i}\}_{i=1}^{T}\text{ with binsize }bin\_size$.
10:     replace $\{v_{i}\}_{i=1}^{T}$ in $C$ with $\{v^{\prime}_{i}\}_{i=1}^{T}$.
11:     go to step 3.
12:  end if
13:  coarsened signal space $\mathbf{V}^{\prime}=\text{unique}(\{v^{\prime}_{i}\}_{i=1}^{T})$
14:  Output: $\hat{\pi}_{\pi^{C}}$, $\mathbf{V}^{\prime}$
Denote the optimal coarsened signal space that satisfies Equation 12 as
[…]
The above algorithm finds the coarsened signal space $\mathbf{V}^{\prime}$ that has the 2-factor approximation of the optimal coarsened signal space $\mathbf{V}^{*}$, $\lvert\mathbf{V}^{\prime}\rvert\geq\frac{1}{2}\lvert\mathbf{V}^{*}\rvert$, in $O(log(\lvert\mathbf{V}^{*}\rvert))$ time.
B.1 DGP with high-dimensional signals
The above algorithm’s time complexity is linear in the number of dimensions of the signal space. However, in practice, high-dimensional signals, such as images and text that appear in human-AI decision problems can still be computationally infeasible to quantify with the above algorithm. To address this issue, we propose a two-step approach to estimate the DGP from a set of observations. First, we use a reduction model to reduce the dimensionality of the signal space. Second, we use the estimated DGP using the reduced signal space in the above algorithm.
The optimal reduction model and its generated embeddings should fulfill two properties. First, the generated signals should be as informative as possible for the decision task, such that they maintain the maximum information value contained in the original signals. Second, the derived DGP that is implied by the embedding signals should be feasible to estimate from the observed dataset. We quantify the above objectives by a similar algorithm based on Train/Test split.
Appendix C Observational Study on the Effect of ACIV on Human-AI Team Performance
In this section, we present the results of an observational study on the effect of $\operatorname{ACIV}$ on human-AI team performance using the dataset from Vodrahalli et al. [2022].
Data description.
We experiment with the publicly available Human-AI Interactions dataset [Vodrahalli et al., 2022]. The dataset comprises 34,783 unique predictions from 1,088 different human participants on four different binary prediction tasks (“Art”, “Sarcasm”, “Cities” and “Census”). In each of the tasks, human participants provide confidence values about their predictions before ($d^{H}$) and after ($d^{H+AI}$) receiving AI advice from a classifier in the form of the classifier’s confidence values ($d^{AI}$). $d^{H}$, $d^{H+AI}$ and $d^{AI}$ are in the range of $[0,1]$.
Results.
We present the $\operatorname{ACIV}$ of the AI confidence values ($d^{AI}$) over human decisions ($d^{H}$) measured by the payoff function as the mean squared error (MSE) in Figure 2, and show the reduction of the mean squared error (MSE) of the human-AI team over the human-alone baseline in Table 1. We find that between these four tasks, the $\operatorname{ACIV}$ of the AI over human decisions predicts the improvement of the human-AI team performance over the human-alone baseline.
Figure 2: The $\operatorname{ACIV}$ of the AI confidence values over human decisions for different tasks. The error bars are 95% confidence intervals.
Table 1: The $\operatorname{ACIV}$ of the AI confidence values over human decisions and the reduction of mean squared error (MSE) of the human-AI team over the human-alone baseline for different tasks.
Task Human-complementary Info Human MSE Human+AI MSE Reduction in MSE
Art 0.1772 [0.1747, 0.1798] 0.243 0.1847 0.0583
Cities 0.0919 [0.0860, 0.0978] 0.1955 0.1615 0.034
Sarcasm 0.1625 [0.1587, 0.1662] 0.2137 0.1767 0.037
Census 0.1147 [0.1077, 0.1217] 0.1997 0.1861 0.0136
Appendix D Demonstration I: Model Comparison on Chest Radiograph Diagnosis
We apply our framework to a well-known cardiac dysfunction diagnosis task [Rajpurkar et al., 2018, Tang et al., 2020, Shreekumar, 2025]. We demonstrate how our framework can be used in model evaluation for analyzing how much complementary information value a set of possible AI models offers to the radiology reports written by experts.
Figure 3: Information value of all deep-learning models calculated under our framework. The first row represents the human-alone decisions (without considering any AI predictions as additional signals). The other rows are the combinations of the human-alone decisions and the AI predictions from different pre-trained models. We list the AI predictions alone to show the AI-complementary information value offered by human decisions.
D.1 Data and Model
We use data from the MIMIC dataset [Goldberger et al., 2000], which contains anonymized electronic health records from Beth Israel Deaconess Medical Center (BIDMC), a large teaching hospital in Boston, Massachusetts, affiliated with Harvard Medical School. Specifically, we utilize chest x-ray images and radiology reports from the MIMIC-CXR database [Johnson et al., 2019] merged with patient and visit information from the broader MIMIC-IV database [Johnson et al., 2023]. The payoff-related state, cardiac dysfunction $\omega\in\{0,1\}$, is coded based on two common tests, the NT-proBNP and the troponin, using the age-specific cutoffs from Mueller et al. [2019] and Heidenreich et al. [2022]. We label the radiology reports by a rule-based tool [Irvin et al., 2019] and use the labels as the human decisions (without AI assistance) in the diagnosis task to solve the problem of computational feasibility with high-dimensional textual reports. The labels are represented by the symptoms as positive, negative, or uncertain, i.e., $d\in\{+,?,-\}$. We fine-tuned five deep-learning models on the cardiac dysfunction diagnosis task, VisionTransformer [Alexey, 2020], SwinTransformer [Liu et al., 2021], ResNet [He et al., 2016], Inception-v3 [Szegedy et al., 2016], and DenseNet [Huang et al., 2017]. Our training set contains 12,228 images, and the validation set contains 6,115 images. On a hold-out test set with 12,229 images, the AUC achieved by the five models is: DenseNet with $0.77$, Inception v3 with $0.76$, ResNet with $0.77$, SwinTransformer with $0.78$, and VisionTransformer with $0.80$.
We consider Brier score, a.k.a., quadratic score, as the payoff function: $S(\omega,d)=1-(\omega-d)^{2}$. The scale of the quadratic score is $[0,1]$ and a random guess ($d\sim\text{Bernoulli}(0.5)$) achieves $0.75$ payoff. We use the quadratic score instead of the mean absolute error that is usually used in cardiac dysfunction diagnosis task because the quadratic score is a proper scoring rule where truthfully reporting the belief maximizes the payoff [^5] We prefer a proper scoring rule so that the rational decision-maker’s strategy is to reveal their true belief, ensuring that the signal’s information value accurately reflects its role in forming beliefs.. We also conduct a robust analysis considering various V-shaped payoff functions with different kinks on a discretized grid of $[0,1]$ with a step of $0.01$. We use the hold-out test set to estimate the data-generating process, which defines the joint distribution of state, human decisions, and AI models’ predictions.
We construct the scale of performances by a no-information bound and a full-information bound. The no-information bound is $\mathrm{R}^{\pi,S}(\emptyset)$, the baseline as we define the information value. The full-information bound is defined as the expected payoff of a rational DM who has access to all signals, human label from radiology report and predictions from five AI models. We do not consider the radiology images as a basic signal in this task because of the problem of computational feasibility with high-dimensional signals. We use the five vision models’ predictions as embeddings of the original images. See more detail on the algorithm of choosing the optimal embeddings in LABEL:app:high-dimensional.
D.2 Results
Can the AI models complement human judgment?
We first analyze the agent-complementary information values in Figure 3, using the Brier score as the payoff function. We find that all AI models provide complementary information value to the aforementioned human judgment. As shown in Figure 3 (comparison between $\mathrm{R}^{\pi,S}(D^{\text{Human}}\cup D^{\text{AI}})$ and $\mathrm{R}^{\pi,S}(D^{\text{Human}})$), all AI models capture at least $20\%$ of the total available information value (across all AI model and human decisions) that is not exploited by human decisions. This motivates deploying an AI to assist humans in this scenario.
In the other direction, the human decisions also provide complementary information to all AI models, comparing $\mathrm{R}^{\pi,S}(D^{\text{Human}})$ with $\mathrm{R}^{\pi,S}(D^{\text{AI}})$ in Figure 3. This observation might inspire, for example, further investigation of the information the humans can access to that is not represented in AI training data.
Which AI model offers the most decision-relevant information over human judgments?
Figure 3 shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models. We assess the stability of VisionTransformer’s superiority over the other AI models across many possible losses to test if there is a Blackwell ordering of models. By Proposition 3.7, we test the payoff of models on all V-shaped scoring rules, shown in Figure 6. Across all the V-shaped payoff functions, we find that VisionTransformer is Blackwell more informative and Inception v3 is Blackwell less informative than all other models. The VisionTransformer achieves a higher information value on all V-shaped scoring rules, implying a higher information value on all decision problems.
Appendix E Demonstration II: Behavioral Analysis on Deepfake Detection
We apply our framework to analyze a deepfake video detection task [Dolhansky et al., 2020], where participants are asked to judge whether a video was created by generative AI, including with the assistance of an AI model.
Figure 4: Information value calculated under our framework in the information model defined by the experiment of Groh et al. [2022]. Basic signals include the seven video level features and three types of agent decisions. The baseline on the left represents the expected payoff given no information, i.e., $\mathrm{R}^{\pi,S}(\emptyset)$, and the benchmark on the right represents the expected payoff given all available information, i.e., $\mathrm{R}^{\pi,S}(\{\Sigma_{1},\ldots,\Sigma_{n},D^{\text{Human}},D^{\text{Human-AI}},D^{\text{AI}}\})$. All the payoffs are calculated by $\mathrm{R}^{\pi,S}(\cdot)$, where $\cdot$ is the signal on the y-axis.
E.1 Data and Model
We define the information model on the experiment data of Groh et al. [2022]. Non-expert participants (n=5,524) were recruited through Prolific and asked to examine the videos. They reported their decisions in two rounds. They first reviewed the video and reported an initial decision ($D^{\text{Human}}$) without access to the AI model. Then, in a second round, they were provided with the recommendation ($D^{\text{AI}}$) of a multitask cascaded convolutional neural network [Zhang et al., 2016], with estimated $65\%$ accuracy on holdout data, and chose whether to change their initial decision. This produced a final decision ($D^{\text{Human-AI}}$). Both decisions were elicited as a percentage indicating how confident the participant was that the video was a deepfake, measured in $1\%$ increments: $d\in\{0\%,1\%,\ldots,100\%\}$. We round the predictions from the AI model to the same 100-scale probability scale available to study participants.
We use the Brier score as the payoff function, with the binary payoff-related state: $\omega\in\{0,1\}=\{\text{genuine},\text{fake}\}$. This choice differs from the mean absolute error used by Groh et al. [2022], but again we use the quadratic score because it is a proper scoring rule where truthfully reporting the belief maximizes the score.
We identify a set of features that were implicitly available to all three agents (human, AI, and human-AI). Because the video signal is high dimensional, we make use of seven video-level features using manually coded labels by Groh et al. [2022]: graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin, all of which are labeled as binary indictors. These are the basic signals in our framework. We estimate the data-generating process $\pi$ using the realizations of signals, state, first-round human decisions, AI predictions, and second-round human-AI decisions. The no-information bound is the same as Appendix D while the full-information bound is defined as the expected payoff of a rational DM who has access to all video-level features and three agents’ decisions.
E.2 Results
How much decision-relevant information do agents’ decisions offer?
We first compare the information value of the AI predictions to that of the human decisions in the first round (without AI assistance). Figure 4(a) shows that AI predictions provide about $65\%$ of the total possible information value over the no-information baseline, while human decisions only provide about $15\%$. Because the no information baseline, $0.75$, is equivalent to a random guess drawn from $\text{Bernoulli}(0.5)$, human decisions are only weakly informative for the problem.
We next consider the human-AI decisions. Given that the AI predictions contain a significant portion of the total possible information value, we might hope that when participants have access to the AI predictions, their performance will be close to the full information baseline. However, the information value of the human-AI decisions only achieves a small proportion of the total possible information value ($30\%$). This is consistent with the findings of Guo et al. [2024] that humans are bad at distinguishing when AI predictions are correct.
How much additional decision-relevant information do the available features offer over agents’ decisions?
To understand what information might improve human decisions, we assess the $\operatorname{ACIV}$s of different signals over different agents. This describes the additional information value in the signal after conditioning on the existing information in the agents’ decisions. As shown on the fifth row in Figure 4, the presence of a flickering face offers larger $\operatorname{ACIV}$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information. Meanwhile, as shown on the fourth row in Figure 4, the presence of an individual with dark skin offers larger $\operatorname{ACIV}$ over AI predictions than over human decisions, suggesting that humans make greater use of this information. This suggests that the AI and human rely on differing information to make their initial predictions, where the AI relies more on information associated with the presence of a flickering face while human participants rely more on information associated with the presence of an individual with dark skin.
By comparing the $\operatorname{ACIV}$s of different signals over human decisions and human-AI decisions, we also find that simply displaying AI predictions to humans did not lead to the AI-assisted humans exploiting the observed signals in their decisions. As shown in Figure 4, with the assistance of AI, the $\operatorname{ACIV}$s of all signals over the human-AI teams’ decisions do not change much compared to the $\operatorname{ACIV}$s over human decisions, with the exception of a slight improvement in the presence of a flickering face. This finding further confirms the hypothesis that humans are simply relying on AI predictions without processing the information contained in them.
Appendix F Demonstration III: Information-based Local Explanation on Recidivism Prediction
We apply our framework to a recidivism prediction task, where the decision-maker decides whether to release a defendant. We demonstrate using the framework to augment SHAP.
Figure 5: (A) Summary of SHAP explanation and (B) summary of ILIV-SHAP explanation, where each point represents an instance from the test set. (C-F) Waterfall plots of applying SHAP and ILIV-SHAP on instance 4 and instance 18. The Waterfall plot starts from the prior expected prediction (or ILIV of the prior expected prediction), and then adds features one at a time until reaching the current model output (or ILIV of the current model output). See further comparisons between SHAP and ILIV-SHAP in Appendix H.
F.1 Data and Model
We use the COMPAS dataset [Angwin et al., 2022], which contains 11,758 defendants with associated features capturing demographics, information about their current charges, and their prior criminal history. We merge the dataset with the experimental data from Lin et al. [2020] to obtain human decisions, in the form of recidivism decisions on a subset of instances from Angwin et al. [2022]’s experiment on laypeople. Human decisions were elicited on a 30-point probability scale. Merging the two datasets produces 9,269 instances (defendants). For the AI model, we trained an extreme gradient boosting (XGBoost) model [Chen and Guestrin, 2016] on a training set with 6,488 instances, achieving an AUC of $0.84$ on a hold-out test set with 2,781 instances. We round the model predictions to the same 30-point probability scale available to study participants in Lin et al. [2020].
We use the Brier score as the payoff function, which is commonly used in previous studies on recidivism [Green and Chen, 2019a, Lin et al., 2020, Fogliato et al., 2021]. The binary payoff-related state $\omega$ indicates whether the defendant gets rearrested within two years. We use the features of the defendant contained in the COMPAS dataset as the signals in our demonstration: demographic features (age, sex, and race), information about the current charge (type of offense and whether it is a misdemeanor or a felony), prior criminal history (e.g., past number of arrests and charges for several offense categories), and the predicted score from the COMPAS system [^6] https://doc.wi.gov/Pages/AboutDOC/COMPAS.aspx. We take the empirical distribution of the hold-out test set as the data-generating process, which defines the joint distribution of state, signals, human decisions, and AI predictions.
F.2 Results
SHAP explanations do not communicate the complementary information of AI predictions.
Specifically, the changes in information value of AI predictions caused by marginalizing out features, i.e., ILIV-SHAP scores, are not predictable from the changes in AI predictions, i.e., SHAP scores, and the feature values. Figure 5 (A) and (B) compare the distribution of feature importance scores from SHAP to those from ILIV-SHAP. The scores of prior_counts (the number of prior criminal charges), age, and decile_score (the score from the COMPAS system), which are three features with the largest scores in both SHAP and ILIV-SHAP, show different correlations with the feature values in SHAP and ILIV-SHAP. For instance, in Figure 5(A), most of the low values (blue dots) of prior_counts have a negative SHAP score, and most of the high values (red dots) have a positive SHAP score. This indicates that when the value of prior_counts is low, marginalizing out prior_counts from the features leads to a higher expected model prediction than not marginalizing out, and vice versa. However, this correlation does not hold for ILIV-SHAP. In Figure 5(B), the high values of prior_counts appear on both the left side (below zero) and the right side (above zero), indicating having low ILIV-SHAP scores and high ILIV-SHAP scores at the same time. This mismatch between the distributions of SHAP scores and ILIV-SHAP scores implies that the impact of features on the information value cannot be directly inferred (at least monotonically) from their impact on model output and their values.
ILIV-SHAP can augment SHAP on communicating the complementary information of AI predictions.
We argue that ILIV-SHAP can supplement saliency-based explanations to provide users with a sense of how the AI predictions can help improve their decisions. With ILIV-SHAP, users can check whether the prediction contains decision-relevant information by looking at the total $\operatorname{ILIV}$ of the AI prediction, which indicates the maximum improvement in expected payoff that could be achieved by incorporating this prediction. Figure 5(D) and (F) show two different cases, where the AI predictions contain substantial human-complementary information value in (D) (shown by the distance between the tail and the head of the Waterfall plot) but very little information in (F).
ILIV-SHAP also tells the user whether an AI model’s prediction encodes useful complementary information from a particular feature. In Figure 5(E), the defendant’s prior count record (prior_counts=1) ranks second among features that impact the AI prediction ($-0.139$). Upon seeing this, users might be interested in investigating how the AI uses prior_counts=1. However, Figure 5(F) shows that prior_counts=1 contributes marginally to the decision-relevant information in AI predictions (about $+0.007$). Thus, with ILIV-SHAP, users can see that even if prior_counts=1 changes the AI prediction significantly, the AI does not predict more accurately as a result of it.
Appendix G Robustness Analysis in Demonstration I
Figure 6: Robust analysis for experiment I on all V-shaped payoff functions. The kink $\mu$ is shown on the x-axis. Each subplot displays the difference between the $\operatorname{ACIV}$ on the row model over human decisions and the $\operatorname{ACIV}$ on the column model over human decisions. A positive value (colored in blue) at $\mu$ indicates the model on the row contains more informative than the model on the column under the evaluation of V-shaped scoring rule with kink $\mu$. The subplots are symmetric along the diagonal, e.g., $(1,2)$ subplot and $(2,1)$ subplot display the same distribution with opposite signs.
Appendix H More examples of ILIV-SHAP
(a) SHAP on instance 0
(b) ILIV-SHAP on instance 0
(a) SHAP on instance 1
(b) ILIV-SHAP on instance 1
(a) SHAP on instance 2
(b) ILIV-SHAP on instance 2
Appendix I Screenshots of the Experiment
Figure 10: Screenshot of the human-alone trials in the first round of the experiment.
Figure 11: Screenshot of the no explanation condition in the second round of the experiment.
Figure 12: Screenshot of the SHAP condition of AI1 in the second round of the experiment.
Figure 13: Screenshot of the SHAP condition in the second round of the experiment.
Figure 14: Screenshot of the ILIV-SHAP + SHAP condition of AI1 in the second round of the experiment.
Figure 15: Screenshot of the ILIV-SHAP + SHAP condition of AI2 in the second round of the experiment.